Lavage Rapide is a Canadian company that owns and operates a large automatic carwash facility near Montreal. The following table provides data concerning the company’s costs:


Fixed Cost
per Month
Cost per
Car Washed
  Cleaning supplies         $ 0.80     
  Electricity   $ 1,100      $ 0.07     
  Maintenance         $ 0.20     
  Wages and salaries   $ 5,000      $ 0.40     
  Depreciation   $ 8,200           
  Rent   $ 2,000           
  Administrative expenses   $ 1,400      $ 0.04     


For example, electricity costs are $1,100 per month plus $0.07 per car washed. The company actually washed 8,400 cars in August and expects to collect an average of $6.10 per car washed.


Required:

Complete the company’s flexible budget for August.
Answer #1
Flexible budget
Actual Cars washed 8400
Revenue 51240 =8400*6.10
Expenses:
Cleaning supplies 6720 =8400*0.80
Electricity 1688 =1100+(8400*0.07)
Maintenance 1680 =8400*0.20
Wages and salaries 8360 =5000+(8400*0.40)
Depreciation 8200
Rent 2000
Administrative expenses 1736 =1400+(8400*0.04)
Total expenses 30384
Net operating income 20856